# Meromorphic functions whose action on their Julia sets is Non-Ergodic

### v2

@@ 1. Introduction
In the same vein, Bock [Bock] proved a dichotomy theorem for transcendental functions: either the Julia set is the whole sphere and for any set $A$ of positive measure, the orbits of almost all points in ${\mathbb{C}}$ have infinitely many iterates that land in $A$, or the Julia set is not the whole sphere, and almost every point in the Julia set is attracted to the post-singular set. Unlike the rational case, though, when the Julia set is the whole sphere, it is not known when such functions are ergodic. Misiurewicz [Mis] has shown that when $f(z)=e^{z}$, the Julia set is the whole sphere, but Lyubich [Lyu2], following earlier work [GGS, Dev], has shown that its action is not ergodic. More examples of entire maps that are ergodic can be found in [WZ, CW]. Skorulski,  [S, S1], answered the ergodicity question for meromorphic maps of the form $f(z)=\lambda\tan z$ and $f(z)=\frac{ae^{z^{p}}+be^{-z^{p}}}{ce^{z^{p}}+de^{z^{-p}}}$ that have two asymptotic values.
A more general sort of affirmative result can be found in [KK2]: If the orbits of the singular values either land on or accumulate on a compact repelling set, and if the $f$ satisfies an additional growth rate condition, its Julia set is the whole sphere and it is ergodic.
In this paper, we study the ergodicity question for meromorphic functions that have some finite number of asymptotic values and no critical values, the so-called Nevanlinna functions. The dynamical properties of these functions were first investigated in [DK], where it was shown that they share many of the dynamical properties of rational maps. In [CJK4], the authors studied the ergodicity question for a subfamily. We proved:
@@
If $f$ is a Nevanlinna function whose asymptotic values all have orbits that land on infinity, then the Julia set is the Riemann sphere, and the action of $f$ is not ergodic there.
@@
For almost every point $z\in\mathbb{C}$,
@@
Theorem A, together with Main Theorems 1,2 and 3, give a full answer to ergodicity questions for Nevanlinna functions whose Julia set is the whole sphere and whose asymptotic values either have finite orbits or land on a compact repeller. The questions of whether the conclusions of all three theorems hold for Nevanlinna functions with asymptotic values whose orbits are infinite and whose accumulation sets are unbounded are still open. Another open question is whether if all the asymptotic values of $f$ have finite orbits, there exists a unique $\sigma$-finite $f$-invariant measure absolutely continuous with respect to Lebesgue measure.
The paper is organized as follows. The first part of Section 2 contains basic definitions and standard theorems. The remainder of that section is devoted to a detailed review of the properties of Nevanlinna functions. In particular, their behavior in a neighborhood of infinity is described by using auxiliary variables, decomposing the map into several factors and finding basic estimates for the factors. Section 3 contains the heart of the proof of Main Theorem 1. Distinct invariant wandering sets are constructed and their Lebesgue measure is shown to be positive proving the function is not ergodic. The final section contains the proofs of Main Theorems 2 and 3.
@@ 2.1. Definitions and standard theorems
Let $f:D(z_{0},r)\to\mathbb{C}$ be a univalent function, and $\eta<1$. Then, for any $A,B\subset D(0,\eta r)$,
@@ 2.2. Nevanlinna functions
In this paper, we continue our study of the properties of the dynamical systems these functions generate. Given a Nevanlinna function, we can define the orbit $\{f^{n}{z}\}$ for any point $z\in{\mathbb{C}}$. In general, these orbits are infinite, but if, for some $m$, $f^{m}(z)=\infty$, the orbit is finite. The set of points with finite orbit is the set of prepoles.
The standard defintions of stable (Fatou) and unstable (Julia) behavior for rational maps carries over to points with infinite orbits; the prepoles are unstable. In [DK], it is proved that the classification of stable behavior for Nevanlinna maps is the same as that for rational maps. In particular, if all the singular points are unstable, the Julia set is the whole sphere.
@@
If $f$ is a Nevanlinna function with $N$ asymptotic values $\lambda_{1},\ldots,\lambda_{N}$, and if for some $0\leq k<N$, $\lambda_{i},i=1,\ldots,k$ are prepoles, and if the $\omega$-limit sets of the remaining $N-k$ asymptotic values are compact repellers, then the Julia set is the Riemann sphere and $f$ is ergodic there.
@@
If $f$ is a Nevanlinna function with $N$ asymptotic values $\lambda_{1},\ldots,\lambda_{N}$, and if ALL of them are prepoles, then $f$ is not ergodic on its Julia set which is the Riemann sphere $\widehat{\mathbb{C}}$.
@@ 2.3. The behavior of $f$ in a neighborhood of infinity
Let $f$ be a Nevanlinna function with Schwarzian derivative $S(f)=2P(z)$, where the degree of the polynomial $P(z)$ is $N-2$ and the leading coefficient of $P(z)$ is the constant $a\in\mathbb{C}^{*}$. Denote solutions to the congruence
@@
In each sector $S_{i}$, the function $f$ has a “truncated solution” which grows like $exp(z^{N/2})$. More precisely, let $R>>0$ and $A_{R}=\{z\ :\ |z|>R\}$. Then in $S_{i}$, one can define the auxilliary variable
[…]
@@
For $z\in S_{i}$ and ${\mathcal{Z}}_{i}={\mathcal{Z}}_{i}(z)$, we obtain a function on ${\mathcal{S}}_{i}$ defined by $F({\mathcal{Z}}_{i})=f(z)$. The map $F$ can be approximately expressed as
@@
are mapped by $f$ to punctured neighborhoods of the asymptotic values $\lambda_{i}$ and $\lambda_{i-1}$ of $f$, $\lambda_{i}$ and $\lambda_{i-1}$ are also the asymptotic values of $F$; that is,
@@
By hypothesis, $f^{k_{i}-1}$ and $f^{k_{i-1}-1}$ map $N_{i}$ and $N_{i-1}$ to neighborhoods $P_{i}$ and $P_{i-1}$ of the poles $p_{i}$ and $p_{i-1}$. Since $f$ maps both $P_{i}$ and $P_{i-1}$ to a neighborhood of infinity, the map $g(z)=1/f(z)$ maps both $P_{i}$ and $P_{i-1}$ into a neighborhood of the origin.
Thus
@@
map $D$ into a neighborhood of 00, and fixes the origin.
Figure 2. The decomposition of $h_{i,U}\circ E_{i,U}$ as a map from the auxilliary plane to the dynamic plane
Set $I(z)=1/z$ and define the following maps on $T_{i}$ and $T_{i-1}$ in the sector $S_{i}$ respectively as:
@@
The maps $\varphi_{i,U}$ and $\varphi_{i,L}$ map the respective asymptotic tracts $T_{i}$ and $T_{i-1}$ onto a neighborhood $\Omega$ of $\infty$. See figures 2 and 3. Next, for $z\in T_{i}\cup T_{i-1}\subset S_{i}$, we define the maps:
$\Phi_{i}(z)=\begin{cases}\phi_{i,U}(z),z\in T_{i}\\ \phi_{i,L}(z),z\in T_{i-1}.\end{cases}$
@@
then $h_{i,U}(\xi)=h_{i+1,L}(\xi)$; that is, $\phi_{i,U}(z)=\phi_{i+1,L}(z)$ so that the map $\Phi_{i}$ is well-defined.
[…]
@@ 2.4. Basic calculations
[…]
@@
(4) $|(\arg(\Phi_{j}(z_{2}))-\theta_{j^{\prime}})|\geq\frac{4}{N\alpha_{k}},\text{ or equivalently }(\sin({\mathcal{Z}}_{j^{\prime}}(\Phi_{j}(z_{2})))\geq\frac{2}{\alpha_{k}},$
then $\Im({\mathcal{Z}}_{j^{\prime}}\Phi_{j}(z_{2}))\geq\Im({\mathcal{Z}}_{i^{\prime}}\Phi_{i}(z_{1}))+2\alpha_{k+1}$.
@@
(6) $\begin{split}\Im({\mathcal{Z}}_{j^{\prime}}(\Phi_{j}(z_{2})))&\geq|(\Phi_{j}(z_{2})\sin(\arg({\mathcal{Z}}_{j}(\Phi_{j}(z_{2})))|\\ &\geq\frac{2|a|^{\frac{N}{2}}}{N}(m_{j}(e^{2y_{2}}-2e^{2c}+e^{4c-2y_{2}}))^{\frac{N}{2}}\frac{2}{\alpha_{k}}.\\ \end{split}$
Note that $y_{2}>y_{1}+2\alpha_{k}$, thus for $k$ large enough,
$\Im{\mathcal{Z}}_{j^{\prime}}(\Phi_{j}(z_{2}))\geq\Im{\mathcal{Z}}_{i^{\prime}}(\Phi_{i}(z_{1}))+2\alpha_{k+1}.$
@@
The pull back of these strips
@@
The image $\Phi_{i}(H_{k}^{i})$ is an annular region $A_{i}$ in a neighborhood of infinity of the $z$-plane that overlaps all of the sectors $S_{j}$’s, and in particular, all the asymptotic tracts $T_{j}$, all the critical lines $L_{j}$ and all the wedges $wed_{j}$. See Figure 3. Thus, the maps ${\mathcal{Z}}_{j}$’s are well-defined for all $1\leq j\leq N$. Define the map $\Psi_{i,j}$ on the ${\mathcal{Z}}_{i}$-plane by the functional equation
@@
The maps $\Psi_{i,j}$ are infinite to one. To create regions of injectivity, divide each horizontal strip $Hor_{k}^{i}$ into infinitely rectangles $Rect_{k,n}^{i}$ of width $\pi$; and define $N$ disjoint sub-rectangles $Rect_{j,k,n}^{i}$ in each rectangle $Rect_{k,n}^{i}$ as:
@@
The vertical lines $x=x_{j}^{i}+n\pi$ in $Hor_{k}^{i}$ are essentially mapped to the critical ray $L_{j}$ under the map ${\mathcal{Z}}_{j}^{-1}\circ\Psi_{i,j}$ and they lie in the complements of the $Rect_{j,k,n}^{i}$.
@@
The next lemma states that in the ${\mathcal{Z}}_{i}$-plane, for any point $x+iy$ in these sub-rectangles $Rect_{j,k,n}^{i}$, the inequality (4) in Lemma 2.5 is satisfied.
@@
$E_{U}(R_{j,k})=e^{-2y}e^{2i(x_{j}-\frac{3}{N\alpha_{k}}})\mbox{ and }E_{U}(M_{j})=e^{-2y}e^{2i(x_{j})}.$
They are line segments meeting at 00 and the angle between them satisfies
@@
Next, since $h_{U}:D(0,\rho)\to\mathbb{C}$ is conformal, the images $h_{U}(E_{U}(M_{j}))$ and $h_{U}(E_{U}(R_{j,k}))$ are curves with the initial point 00. Let $\gamma_{M}(t)$ and $\gamma_{R}(t)$ be the lines tangent to these curves at 00, respectively. Then
@@
$I(w)=\frac{1}{r_{0}}e^{-i(\arg h_{U}^{\prime}(0)+2x_{j})}\text{ and }\Phi(z)=I(h_{U}(\xi))=\frac{e^{-i\theta}}{r_{0}}.$
@@
$\arg(\Phi_{i}({\mathcal{Z}}_{i}^{-1}(z)))-\theta_{j}>4/(N\alpha_{k})$. Since ${\mathcal{Z}}$ was an arbitrary point in $R_{j,k}$, we conclude
$\arg(\Psi_{i,j}({\mathcal{Z}}))\geq 2/\alpha.$
@@
Since ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$, by lemma 2.6, $|\arg(\Phi_{i,j}({\mathcal{Z}}))|>2/\alpha_{k}$. As ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$, then $\Im{\mathcal{Z}}_{j}\geq\alpha_{k}+2\alpha_{k-1}$, thus from Lemma 2.5 we have
@@
3. Disjoint wandering sets of positive measure

### v3

@@ 1. Introduction
In the same vein, Bock [Boc] proved a dichotomy theorem for transcendental functions: either the Julia set is the whole sphere and for any set $A$ of positive Lebesgue measure, the orbits of almost all points in $\mathbb{C}$} have infinitely many iterates that land in $A$, or the Julia set is not the whole sphere, and almost every point in the Julia set is attracted to the post-singular set. Unlike the rational case, though, when the Julia set is the whole sphere, it is not known when such functions are ergodic. Misiurewicz [Mis] has shown that when $f(z)=e^{z}$, the Julia set is the whole sphere, but Lyubich [Lyu2], following earlier work [GGS, Dev], has shown that its action is not ergodic. More examples of entire maps that are ergodic can be found in [WZ, CW]. Skorulski,  [S, S1], answered the ergodicity question for meromorphic maps of the form $f(z)=\lambda\tan z$ and $f(z)=\frac{ae^{z^{p}}+be^{-z^{p}}}{ce^{z^{p}}+de^{z^{-p}}}$ that have two asymptotic values.
A more general sort of affirmative result can be found in [KK2]: If the orbits of the singular values either land on or accumulate on a compact repelling set, and if the mapff satisfies an additional growth rate condition, then its Julia set is the whole sphere and it is ergodic.
In this paper, we study the ergodicity question for meromorphic functions whose singular set consists of finitely many asymptotic values and no critical values, the so-called Nevanlinna functions. The dynamical properties of these functions were first investigated in [DK], where it was shown that they share many of the dynamical properties of rational maps. In [CJK4], the authors studied the ergodicity question for a subfamilies of Nevanlinna functions for which some of the asymptotic values have orbits that land on infinity. They proved:
@@
Let $f$ be a Nevanlinna function with $N$ asymptotic values, $\lambda_{1},\ldots,\lambda_{N}$. If all $N$ asymptotic values land on infinity, then the Julia set of $f$ is the Riemann sphere, and the action of $f$ is not ergodic there.
@@
For almost every point $z\in\mathbb{C}$, its $\omega$-limit set satisfies:
@@
Theorem A, together with Main Theorems 1,2 and 3, give a full answer to ergodicity questions for Nevanlinna functions whose Julia set is the whole sphere and whose asymptotic values land either on infinity or on a compact repeller. The questions of whether the conclusions of all three theorems hold for Nevanlinna functions with asymptotic values whose orbits are infinite and whose accumulation sets are unbounded are still open. Another open question is whether there exists a unique $\sigma$-finite $f$-invariant measure absolutely continuous with respect to Lebesgue measure if all the asymptotic values of $f$ have finite orbits.
The paper is organized as follows. The first part of Section 2 contains basic definitions and standard theorems. The remainder of that section is devoted to a detailed review of the properties of Nevanlinna functions. In particular, the asymptotic values divide a neighborhood of infinity into asymptotic tracts on which the function is exponentially contracting. In order to prove the function is non-ergodic, we need to produce disjoint invariant wandering sets in these tracts, each of which has positive Lebesgue measure. This involves estimating the contraction on the asymptotic tracts. In order to make these estimates, we use some classical techniques. We introduce an “auxiliary” variable in each tract which converts it to a half plane in which we can make the estimates. Section 3 contains the heart of the proof of Main Theorem 1. Distinct invariant wandering sets are constructed and their Lebesgue measure is shown to be positive proving the function is not ergodic. The final section contains the proofs of Main Theorems 2 and 3.
@@ 2.1. Definitions and standard theorems
Let $f:D(z_{0},r)\to\mathbb{C}$ be a univalent function, and $\eta<1$. Then, for any $A,B\subset D(z_{0},\eta r)$,
@@ 2.2. Nevanlinna functions
In this paper, we continue our study of the properties of the dynamical systems these functions generate. Given a Nevanlinna function, we can define the orbit $\{f^{n}{z}\}$ for any point $z\in{\mathbb{C}}$. In general, these orbits are infinite, but if, for some $m\geq 0$, $f^{m}(z)=\infty$, the orbit is finite. Such points are called prepoles of order $m$.
The standard defintions of stable (Fatou) and unstable (Julia) behavior for rational maps carry over to points with infinite orbits; the prepoles are unstable. In [DK], it is proved that the classification of stable behavior for Nevanlinna maps is the same as that for rational maps. In particular, if all the singular points are unstable, the Julia set is the whole sphere.
@@
If $f$ is a Nevanlinna function with $N$ asymptotic values $\lambda_{1},\ldots,\lambda_{N}$, and if for some $0\leq k<N$, $\lambda_{i},i=1,\ldots,k$ are prepoles, [^1] If some $\lambda_{i}=\infty$, it is a ”prepole of order 0”. and if the $\omega$-limit sets of the remaining NkN-k asymptotic values are compact repellers, then the Julia set is the Riemann sphere and $f$ is ergodic there.
@@
If $f$ is a Nevanlinna function with $N$ asymptotic values $\lambda_{1},\ldots,\lambda_{N}$, and if ALL of them are prepoles, then its Julia set is the Riemann sphere $\widehat{\mathbb{C}}$ and $f$ is not ergodic.
Before getting into the details we sketch the main ideas in the proof (see also [Lyu1, S]). We construct two disjoint sets $A$ and $B$ such that:
  1. (1)
    Both $A$ and $B$ have positive Lebesgue measure;
  2. (2)
    Both sets are wandering and their orbits are disjoint.
Form the sets $X_{A}=\cup_{n\geq 0}f^{-n}(\cup_{n\geq 1}f^{n}(A))$ and $X_{B}=\cup_{n\geq 0}f^{-n}(\cup_{n\geq 1}f^{n}(B))$. They are both invariant and both have positive Lebesgue measure; this proves that the map is non ergodic.
For the construction of $A$ and $B$, we see that the map $f$ has $N$ asymptotic tracts and $N$ asymptotic values each of which is a prepole. Therefore, each asymptotic tract is expansively mapped over all $N$ asymptotic tracts under some finite number of iterations. We fix an asymptotic tract and choose some quadrilaterals that, under an appropriate change of coordinate are squares. We then show that the Lebesgue measure of the set of points whose orbit always lands on these $N$ asymptotic tract is positive and it wanders to $\infty$; these are ”almost” Markov chains because the map is expanding. If we choose the quadrilaterals so that the distance between them is large enough, since the expansion rate on the tracts are different, their orbits will be disjoint.
@@ 2.3. The behavior of $f$ in a neighborhood of infinity
Let $f$ be a Nevanlinna function with Schwarzian derivative $S(f)=2P(z)$, where the degree of the polynomial $P(z)$ is N2N-2 and the leading coefficient of $P(z)$ is the constant $a\in\mathbb{C}^{*}$. The Schwarzian equation is related to the linear equation $w^{\prime\prime}+P(z)w=0$: that is, if $w_{1},w_{2}$ are linearly independent solutions of this equation, then
$S(\frac{aw_{1}+bw_{2}}{cw_{1}+dw_{2}})=2P(z).$
A special case is $N=2$ where the solutions $w_{1},w_{2}$ can chosen as exponentials whose asymptotic tracts are half planes.
This linear equation is often studied by making changes of variable in the asymptotic tracts in order to estimate the behavior of $f$ there. The new variable turns the asymptotic tract into a half plane and the Liouville transformation turns the Schwarzian equation into an approximation of the equation in the case where $N=2$. The interested reader can find more details in the discussions in [CJK4] and [L]. We follow the notation in those discussions.
Denote solutions to the congruence
@@
In each sector $S_{i}$, the function $f$ has a “truncated solution” which grows like $exp(z^{N/2})$. More precisely, let $R>>0$ and $A_{R}=\{z\ :\ |z|>R\}$. Then in $S_{i}$, one can define the auxiliary variable
${\mathcal{Z}}_{i}(z)=\int^{z}_{Re^{i\theta_{i}}}\sqrt{P(s)}ds=\frac{2\sqrt{a}}{N}z^{N/2}(1+o(1)),z\in S_{i}\cap A_{R}.$
@@
For $z\in S_{i}$ and ${\mathcal{Z}}_{i}={\mathcal{Z}}_{i}(z)$, we use the Liouville transformation to transform the linear equation to one in the variable ${\mathcal{Z}}_{i}$ that has a solution on ${\mathcal{S}}_{i}$ defined by $F({\mathcal{Z}}_{i}(z))=f(z)$. The map $F({\mathcal{Z}}_{i})$ can be approximately expressed as
@@
are mapped by $f$ to punctured neighborhoods of the asymptotic values $\lambda_{i}$ and $\lambda_{i-1}$ of $f$, $\lambda_{i}$ and $\lambda_{i-1}$ are also the asymptotic values of $F$; that is, [^2] Since this holds for each pair of asymptotic values, $\lambda_{i},\lambda_{i-1}$, it is clear both cannot be infinite.
@@
By hypothesis, $f^{k_{i}-1}$ and $f^{k_{i-1}-1}$ map $N_{i}$ and $N_{i-1}$ to neighborhoods $P_{i}$ and $P_{i-1}$ of the poles $p_{i}$ and $p_{i-1}$ and so $f$ maps both $P_{i}$ and $P_{i-1}$ to a neighborhood of infinity. Because it is simpler to estimate maps between finite regions, we introduce a transformation that maps a neighborhood of infinity into a neighborhood of the origin. To this end, let $I(z)=1/z$ and set
$f=I\circ g,\text{where }I=\frac{1}{z},\,g(z)=\frac{1}{f(z)}$
so that the map $g(z)$ maps both $P_{i}$ and $P_{i-1}$ into a neighborhood of the origin. The maps
@@
both fix the origin and so map the disk $D(0,e^{-c})$ into a neighborhood of 0.
Figure 2. The decomposition of $h_{i,U}\circ E_{i,U}$ as a map from the auxiliary plane to the dynamic plane
Next, define the following maps on $T_{i}$ and $T_{i-1}$ in the sector $S_{i}$ respectively as:
@@
The maps $\varphi_{i,U}$ and $\varphi_{i,L}$ are compositions of a Mobius map, a univalent map and the exponential map, and they map the respective asymptotic tracts $T_{i}$ and $T_{i-1}$ onto a neighborhood $\Omega$ of $\infty$. See figures 2 and 3. Next, for $z\in T_{i}\cup T_{i-1}\subset S_{i}$, we define the maps:
$\Phi_{i}(z)=\begin{cases}\varphi_{i,U}(z),z\in T_{i}\\ \varphi_{i,L}(z),z\in T_{i-1}.\end{cases}$
@@
then $h_{i,U}(\xi)=h_{i+1,L}(\xi)$; that is, $\varphi_{i,U}(z)=\phi_{i+1,L}(z)$ so that the map $\Phi_{i}$ is well-defined.
Recall that the fixed constant $c>>0$ is the large constant that defines the sets ${\mathcal{L}}_{i}$ and ${\mathcal{U}}_{i}$. Assume it is chosen large enough so that for all $1\leq i\leq N$,
@@ 2.4. Basic calculations
Choose an $\alpha_{0}>c$, and for integers k>0k>0333We use $k$ here as an index and $k_{i}$ as the order of the prepole $\lambda_{i}$. This should not cause confusion., where $N$ is the number of asymptotic values of $f$, define the sequences of real numbers,
@@
(4) $|(\arg(\Phi_{j}(z_{2}))-\theta_{j^{\prime}})|\geq\frac{4}{N\alpha_{k}},\text{ or equivalently }|\sin({\mathcal{Z}}_{j^{\prime}}(\Phi_{j}(z_{2}))|\geq\frac{2}{\alpha_{k}},$
then $\Im({\mathcal{Z}}_{j^{\prime}}(\Phi_{j}(z_{2})))\geq\Im({\mathcal{Z}}_{i^{\prime}}(\Phi_{i}(z_{1})))+2\alpha_{k+1}$.
@@
[…]
Note that $y_{2}>y_{1}+2\alpha_{k-1}$, and $\alpha_{k}=e^{N\alpha_{k-1}}$, so that
(7) $\begin{split}(m_{j}(e^{2y_{2}}-2e^{2c}+e^{4c-2y_{2}}))^{\frac{N}{2}}\frac{2}{\alpha_{k}}&>2(m_{j}(\frac{e^{2y_{1}+4\alpha_{k-1}}-2e^{2c}+e^{4c-2y_{2}})}{e^{2\alpha_{k-1}}})^{\frac{N}{2}}\\ &=2(m_{j}(e^{2y_{1}+2\alpha_{k-1}}-\frac{2e^{2c}-e^{4c-2y_{2}})}{e^{2\alpha_{k-1}}})^{\frac{N}{2}}\\ &>(m_{i}(e^{2y_{1}}+2e^{2c}+e^{4c-2y_{1}}))^{\frac{N}{2}}+N\alpha_{k+1}/|a|^{N/2}\end{split}$
for $k$ large enough, since $\alpha_{k}\to\infty$ and $y_{1}\geq\alpha_{k}$. Therefore
$\Im({\mathcal{Z}}_{j^{\prime}}((\Phi_{j}(z_{2}))))\geq\Im({\mathcal{Z}}_{i^{\prime}}(\Phi_{i}(z_{1})))+2\alpha_{k+1}.$
@@
The pull-back of these strips
@@
The image $\Phi_{i}(H_{k}^{i})$ is an annular region $A_{i}$ in a neighborhood of infinity of the $z$-plane that, intersects all of the sectors $S_{j}$ for all $j=1,\ldots,N$, and in particular, all the asymptotic tracts $T_{j}$, all the critical lines $L_{j}$ and all the wedges $wed_{j}$. See Figure 3. Thus, the maps ${\mathcal{Z}}_{j}$’s are well-defined for all $1\leq j\leq N$. Define the map $\Psi_{i,j}$ on the ${\mathcal{Z}}_{i}$-plane by the functional equation
@@
The maps $\Psi_{i,j}$ are infinite to one. To create regions of injectivity, divide each horizontal strip $Hor_{k}^{i}$ into infinitely many rectangles $Rect_{k,n}^{i}$ of width $\pi$ and define $N$ disjoint sub-rectangles $Rect_{j,k,n}^{i}$ in each rectangle $Rect_{k,n}^{i}$ as:
@@
Under the map ${\mathcal{Z}}_{j}^{-1}\circ\Psi_{i,j}$, each rectangle $Rect_{k,n}^{i}$ maps injectively onto the annulus in the $z$-plane. The vertical lines $x=x_{j}^{i}+n\pi$ in $Hor_{k}^{i}$ are preimages of the critical rays $L_{j}$. The rectangles $Rect_{j,k,n}^{i}$ in each $Rect_{k,n}^{i}$ injectively map onto the disjoint asymptotic tracts in the annulus.
@@
The next lemma states that in the ${\mathcal{Z}}_{i}$-plane, for any point $x+iy$ in these sub-rectangles $Rect_{j,k,n}^{i}$, the inequality (4) in Lemma 2.5 is satisfied. In particular, the sub-rectangles are mapped away from the critical rays and into an asymptotic tract.
@@
$E_{U}(R_{j,k})=e^{-2y}e^{2i(x_{j}-\frac{3}{N\alpha_{k}}})\mbox{ and }E_{U}(M_{j})=e^{-2y}e^{2i(x_{j})},\ y\geq\alpha_{0}.$
They are line segments meeting at 0 and the angle between them satisfies
@@
Next, since $h_{U}:D(0,\rho)\to\mathbb{C}$ is conformal, the images $h_{U}(E_{U}(M_{j}))$ and $h_{U}(E_{U}(R_{j,k}))$ are curves with the initial point 0. Let $\gamma_{M}(t)$ and $\gamma_{R}(t)$ be the lines tangent to these curves at 0, respectively. Then
@@
$I(w)=\frac{1}{r_{0}}e^{-i(\arg h_{U}^{\prime}(0)+2x_{j})}\text{ and }\Phi_{i}(z)=I(h_{U}(\xi))=\frac{e^{-i\theta}}{r_{0}}.$
@@
$\arg(\Phi_{i}({\mathcal{Z}}_{i}^{-1}(\mathcal{Z})))-\theta_{j}>4/(N\alpha_{k})$. Note that $\mathcal{Z}$} is an arbitrary point in $R_{j,k}$, and by the definition of $\theta_{j}$, $I(\gamma_{M}(t))$, it is mapped to the positive real line under the map $\mathcal{Z}_{j}$. Since $\Psi_{i,j}=\mathcal{Z}_{j}\circ\Phi_{i}$, we conclude
$\arg(\Psi_{i,j}({\mathcal{Z}}))\geq 2/\alpha_{k}.$
@@
Since ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$, by lemma 2.6, $|\arg(\Phi_{i,j}({\mathcal{Z}}))|>2/\alpha_{k}$. Because ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$, $\Im{\mathcal{Z}}_{j}\geq\alpha_{k}+2\alpha_{k-1}$; thus Lemma 2.5 implies
@@
3. Disjoint wandering sets of positive Lebesgue measure